OpenStudy (anonymous):

find the x intercept of the parabola with vertex (-1,-16) and y intercept (0,-15) write your answer in this form (x1,y1), (x2,y2) Please help!!!!

OpenStudy (mathmate):

Start by drawing a sketch to represent the parabola. Since the y-intercept is higher than the vertex, the parabola is concave up.  The vertex (xv, yv) is a convenient way to write the equation of the parabola: Equation for a parabola that is concave up or down (instead of left/right). y-yv = a(x-xv)^2 given (xv,yv)=(-1,-16), we have the equation: y-(-16) = a(x-(-1))^2 y=a(x+1)^2-16 To find the numerical value of a, we substitute the other known point (0,-15) into the equation and solve for a: -15=a(0+1)^2-16 Solve for a= 1, so the equation of the parabola is: y=(x+1)^2-16 To find the x-intercepts, solve for x in y=(x+1)^2-16=0